OpenStudy (babyslapmafro):

Please help, assume that 10 ft lbs of work is required to stretch a spring 1 ft beyond its natural length. What is the spring constant?

OpenStudy (babyslapmafro):

My answer is 10 lb/ft but the answer in the book is 20lb/ft

OpenStudy (babyslapmafro):

w=fd 10ft lbs = f (1ft) F=10lbs F(x)=kx 10lbs=k(1ft) k=10lbs/ft

OpenStudy (anonymous):

Energy(spring) = 1/2 * k * x^2 Plugging in your numbers, you have: 10 = 1/2 * k * (1)^2 Solving for k, k = 20

OpenStudy (anonymous):

The energy it takes to stretch a spring is \[\frac{kx^2}{2}\]. Therefore, \[E = \textrm{10 ft * lbs} = \frac{k * 1^2}{2}\] Rearranging the equation gives k = 20 lb / ft